The Operations and Engineering personnel aboard JOIDES Resolution for Leg 183 were Michael Storms, ODP Operations Manager, and Steven Kittredge, Schlumberger Engineer.


Transit to Site 1135 (Proposed Site KIP-13B)
At 0745 hr on 13 December 1998, the last line was passed ashore and the ship departed Fremantle. At 0812 hr the pilot was away, and the ship was under way at full speed for the first site of Leg 183. For the first 4 days of the transit (north of 44°S), the ship made good time, averaging just under 10 kt. On the 5th day of the transit, the wind speed increased, blowing predominantly out of the south or southwest, and the seas/swells continued to build over the next 5 days. As local weather conditions deteriorated and the opposing Antarctic Circumpolar Current strengthened, the transit speed dropped significantly, varying between 8.8 and 6.9 kt. Winds of 20 kt or greater increased to 40 to 50 kt over this period, with gusts over 50 kt. Swells increased to a maximum of 20 ft with superimposed wind generated seas of up to 12 ft. Water temperature dropped from an initial 23° to 3°C, and the air temperature decreased from 21° to 3°C. During the final 1 1/2 days of the transit (south of 56°S), the wind began blowing out of the north and the seas abated considerably. Transit speed increased accordingly, averaging better than 10 kt for much of this period. Seawater and air temperatures continued to drop, however, to 0.6° and 0°C, respectively, and snow flurries increased in frequency during the last 1-2 days of the transit. The 2093-nmi transit from Fremantle to Site 1135 was accomplished at an average speed of 8.6 kt. The ship arrived on location for the first drill site on 23 December 1998.

Hole 1135A
At 1145 hr on 23 December 1998, the beacon was deployed on the precise differential Global Positioning System (GPS) coordinates for Site 1135. Using a C-4 core bit, Hole 1135A was spudded with the rotary core barrel (RCB) at 2145 hr on 23 December 1998. A good seafloor tag indication was identified by the driller. The depth to the seafloor adjusted to the rig floor datum was 1578.0 m. The 3.5-kHz precision depth recorder (PDR) indicated an adjusted seafloor depth of 1578.4 m.
Continuous wireline coring proceeded in homogeneous white foraminifer-bearing nannofossil ooze interspersed with thin chert layers. Recovery in the upper 190.0 m was just under 60%. Recovery continued to drop, however, as the percentage of chert increased and the ooze graded into chalk. Sepiolite mud sweeps every other core were initiated while cutting Core 49R at 458.9 mbsf. While installing the sinker bars to retrieve Core 49R, the bit jets plugged. Normal circulating pressures were eventually established after working the pipe off-bottom with the sinker bars out. Coring proceeded to a total depth (TD) of 2104.0 m (526.0 mbsf) before the hole collapsed while attempting to recover Core 55R. This was not surprising since the lowermost portion of the hole from 440.0 to 526.0 mbsf contained abundant chert layers. Recovery through this lowermost part of the section was only 11.4%.
After pulling several stands with 50k overpull and 600 A torque, circulation was regained. A wiper trip to bring the top of the 8 1/4-in drill collars to the seafloor was then undertaken. There were no problems during the trip back to bottom until tagging fill at 447.7 mbsf, ~78 m above the original TD. After 2.75 hr of reaming and circulating mud pills, attempts to clean out the hole were abandoned. A decision was made to move to an alternate location where sediment cover on top of basement was estimated to be only ~150 m thick.
During operations at Site 1135, several icebergs were identified on radar and observed visually. At times, four to five icebergs were at distances between 10 and 15 nmi. However, no icebergs posed a threat to the drilling operation.
While the drill string was being recovered, both the primary and back-up positioning beacons were released and recovered. The ship was secured for transit to Site 1136, a recently approved alternate ~17 nmi to the east. At 2145 hr on 26 December 1998, we got under way in dynamic positioning (DP) mode for Site 1136.

Leg 183 Operations - Site 1136
Leg 183 Table of Contents